# Theory of Mobility Rings in Non-Hermitian Systems

### v1

@@ Abstract
The Hamiltonian for Eq. (14) can be obtained by successive transformations with its exact self-duality relation $E_{c}=(\pm 2-i\lambda)/b$ Sup . We obtain the self-duality points here also by a similar method, which we call hidden duality Sup ; MGoncalves2022 ; MGoncalves2023a ; MGoncalves2023b . The irrational number $\alpha$ can be approximated by the Fibonacci number $F_{m}/F_{m+1}$, knowing only that $\alpha=(\sqrt{5}-1)/2$ for $m\rightarrow\infty$. For the lowest order approximation, $\alpha=1$, the eigenequation for the Hamiltonian can be written as
@@
By the Fourier transform, we can get its eigenvalue as
@@
for $1-b\cos(\theta)\neq 0$, which is self-duality under $\theta\leftrightarrow k$ if $E_{c}=(2-i\lambda)/b$. Defining $\theta^{\prime}=\theta+\pi$, we get $E_{c}=-(2+i\lambda)/b$. This result is the same as the exact self-duality relation. In contrast, the LE obtained by Avila’s global theory is Sup
@@
Figure 4: The fractal dimension $\Gamma$ of all eigenstates in the complex plane. The dashed circle (the two red triangle) denotes the LE given by Eq. (17) (the two self-duality points obtained from the self-duality relation). We fix $L=610$, $\lambda=1$, $b=0.8$, and $\theta=0$.
It can be seen that the MR of the non-Hermitian GPD model is elliptic in the complex plane compared to the mosaic model. We show in Fig. 4 the fractal dimension $\Gamma$ of all eigenstates in the complex plane for $b=0.8$ and $\lambda=1$ (The remaining parameters are detailed in the supplementary material Sup ). The eigenstates inside the MR are all extended eigenstates with $\gamma=0$, and outside the MR are localized states with $\gamma>0$. In other words, MR perfectly separates the domains of the extended and localized states. In contrast, the two self-duality points can only characterize the range of real energies of the extended state under $Im(E)=\lambda/b$. This result reveals that the MEs under non-Hermitian are not well characterized through the self-duality relation, especially for systems without $PT$-symmetry.
Conclusion.—We investigate the MEs of non-Hermitian systems. Due to the existence of complex eigenvalues, we introduce the notion of a “mobility ring” as an extension of the ME in non-Hermitian systems, where the ME is no longer a critical point on the real axis but a closed circle in the complex plane, characterizing extended states inside the ring and localized states outside. By testing different cases, we show that the MR theory is universally applicable to non-Hermitian systems with and without the protection of $PT$-symmetry. Furthermore, multiple pairs of MEs in Hermitian systems will form multiple MRs in non-Hermitian systems. Finally, we compare the results of mobility rings with that of dual transformations, and find that although the self-dual method can give the interval of real eigenvalues corresponding to the extended states, it can not fully display the mobility edge information in the complex plane. The mobility ring theory proposed here is a non-Hermitian counterpart of the mobility edge theory, and we hope this work will bring benefits to the research of related fields.
Note added: In completing this manuscript, we note that a recent preprint entitled “Non-Hermitian butterfly spectra in a family of quasiperiodic lattices LWang2024 ”, similarly investigate the mobility edges in non-Hermitian system. Their article is discussed from the perspective of butterfly spectrum. In this paper, in addition to studying the spectrum, we also discuss the properties of multiple mobility rings and analyze the limitations of self-duality relation in non-Hermitian systems.
Acknowledgements.—This work was supported by the National Key Research and Development Program of China (Grant No.2022YFA1405300), the National Natural Science Foundation of China (Grant No.12074180), and the Guangdong Basic and Applied Basic Research Foundation (Grants No.2021A1515012350).
@@
“Theory of Mobility Rings in Non-Hermitian Systems”
@@
If the energy $E$ is localized for the eigenstates, then its localized length $\Lambda=1/\gamma$ and the wave function of the eigenstate $\psi$ can be given by
$|\psi|\propto|\psi|_{\max}e^{-\gamma|j-j_{c}|},$ (S1)
where $|\psi|_{\max}$ is the maximum value of $|\psi|$ at the center site $j_{c}$ of localization. In Fig. 3(a) we show the fractal dimensions $\Gamma$ of all eigenstates in the complex plane for $\delta=1+i$, $h=1$, and plot the contours of the LE via Eq. (9). One can observe that MR does not depend on whether the system has $PT$-symmetry or not. The extension under non-Hermitian conditions originates from LE in the complex plane. The region with $\gamma=0$ forms one valley in the complex plane (inside the MR), and the eigenstate $\Gamma\rightarrow 1$ in this region is the extended state. As the MR gradually moves outward, the smaller the $\Gamma$, the more localized the corresponding eigenstate. For eigenstates within the MR, as shown in Fig. 3(b), the wavefunction can extend to all lattices. For localized states with $\gamma>0$, as shown in Fig. 3(c)(d), we present the distributions of the eigenstates for $\gamma=0.222$ and $\gamma=0.950$ over all lattices, respectively, and compare them with the fitted Eq. (S1). It is evident that Eq. (S1) effectively characterizes the localization features of the eigenstates, and the larger the LE, the more localized it is.
@@
$\eta\chi_{p}^{-1}(\omega_{0},0)e^{p\theta}u_{p}=G\sum_{r}e^{-\omega|p-r|}e^{r\theta}u_{r}$ (S8)
@@
$\eta\chi_{m}^{-1}(\omega,q\pi)v_{m}=G\sum_{r}e^{-\omega_{0}|m-r|}v_{r}$ (S9)
In the last step we define $z_{k}=\sum_{m}e^{im(2\pi\alpha k+\theta)}v_{m}$. We multiply Eq. (S9) by $e^{im(2\pi\alpha j+\theta)}$ and sum over $m$ and can obtain
@@
$H(\phi,K)=\begin{pmatrix}V(\phi/n_{2})&1&\dots&e^{iK}\\ 1&V(2\phi n_{1}/n_{2}+\phi/n_{2})&\dots&0\\ \vdots&&\ddots&\vdots\\ e^{-iK}&\dots&1&V(2\phi n_{1}(n_{2}-1)/n_{2})+\phi/n_{2})\\ \end{pmatrix}$ (S12)
@@
This can be clearly seen for $\lambda=1$ at $\phi\leftrightarrow K$ and $\lambda=-1$ at $\phi^{\prime}\leftrightarrow K$ the eigenvalue is invariant. Putting the two conditions together, we get the self-duality condition $|\lambda|=1$.
@@
By $\phi\leftrightarrow K$ and $\phi^{\prime}\leftrightarrow K$, we can obtain its self-duality point as $E=(\pm 2-\lambda)/b$ and this result is in agreement with Ref. SGaneshan2015 . In addition, this method supports solving models that do not have self-duality.
@@
Further, we contrast this by Avila’s global theory. we can calculate the LE from the transfer matrix
@@
$\gamma_{B}(E,\epsilon)=\epsilon+\left|\frac{Eb\pm i\lambda+\sqrt{(Eb+i\lambda)^{2}-4b^{2}}}{4}\right|$ (S21)
@@
$\gamma(E,0)=\mathrm{max}\left\{\ln\left|\frac{Eb\pm i\lambda+\sqrt{(Eb+i\lambda)^{2}-4b^{2}}}{2(1+\sqrt{1-b^{2}})}\right|,0\right\}.$ (S22)

### v3

@@ Abstract
where $\lambda$ denotes the quasiperiodic intensity, $\alpha$ is the quasiperiodic parameter, $b$ is a deformation parameter, and $\theta$ denotes a global phase. We set $\alpha=(\sqrt{5}-1)/2$, $|b|<1$. It is clear that the Hamiltonian with the potential Eq. (14) does not have $PT$-symmetry and when $\lambda>0$ the system will have a complex spectrum. The corresponding Hamiltonian can be obtained by successive transformations with its exact self-duality relation $E_{c}=(\pm 2-i\lambda)/b$ Sup . One can obtain the self-duality points by the hidden duality method Sup ; MGoncalves2022 ; MGoncalves2023a ; MGoncalves2023b . The irrational number $\alpha$ can be approximated by the Fibonacci number $F_{m}/F_{m+1}$, i.e., $\alpha=(\sqrt{5}-1)/2$ for $m\rightarrow\infty$. For the lowest order approximation, $\alpha=1$, the eigenequation for the Hamiltonian reads
@@
By the Fourier transform, one can get its eigenvalue as
@@
for $1-b\cos(\theta)\neq 0$, which is self-duality under $\theta\leftrightarrow k$ if $E_{c,1}=(2-i\lambda)/b$. Defining $\theta^{\prime}=\theta+\pi$, we get $E_{c,2}=-(2+i\lambda)/b$. This result is the same as the exact self-duality relation. In contrast, the LE obtained by Avila’s global theory is Sup
@@
Figure 4: $\Gamma$ in the complex plane. The dashed circle denotes the LE given by Eq. (17), whereas the two red triangles arecself-duality points obtained from the self-duality relation. Throughout, $L=610$, $\lambda=1$, $b=0.8$ and $\theta=0$.
Fig. 4 shows the fractal dimensions in the complex plane, where the black dashed ring is MR obtained by Avila’s global theory (the elliptic equation), and the red triangles are the energy critical points obtained by the self-duality relation. MR perfectly divides the extended and the localized regions, and shows how the extended and the localized states distribute in the complex plane, i.e., the extended (localized) states are inside (outside) MR. However, the self-duality relation can merely obtain two special critical points on MR, namely, $Im(E)=\lambda/b$. This result follows from the existence of an infinite number of self-dual points in non-Hermitian systems. That is to say, the self-duality relation is not comprehensive in studying the MEs in the non-Hermitian system.
Conclusion.—In conclusion, MEs in the non-Hermitian system are studied. Since they exhibit a ring structure in the complex plane, we call them “mobility ring”. Further, we analyze MR in various cases and find that MR is universal for the non-Hermitian system, no matter the system has $PT$-symmetry or not. In addition, by analogy with multiple MEs in the Hermitian system, we investigate multiple MRs in complex plane, and give the expression of the maximum number of MRs in the mosaic model. Finally, by comparing the results given by Arvila’s global theory and self-duality relation, we find that self-duality relation has limitations in studying the critical energy in the complex plane. Generally, non-Hermitian systems have complex spectra, which means that MR studied in this paper is universal in the complex plane. Hopefully our findings will bring benefits to the study of ME theory and non-Hermitian physics.
Note added: In completing this manuscript, we note that a recent preprint entitled “Non-Hermitian butterfly spectra in a family of quasiperiodic lattices”, similarly investigates the mobility edges in non-Hermitian system LWang2024 . Their article is from the perspective of butterfly spectrum. In this paper, in addition to studying the spectrum, we further discuss the properties of multiple mobility rings and clarify the limitations of self-duality method in non-Hermitian cases.
Acknowledgements.—We thank Qi Zhou for his insightful suggestions. This work was supported by the National Key Research and Development Program of China (Grant No.2022YFA1405300), the National Natural Science Foundation of China (Grant No.12074180), and the Guangdong Basic and Applied Basic Research Foundation (Grants No.2021A1515012350).
@@
“Ring Structure in the Complex Plane: A Fingerprint of non-Hermitian Mobility Edge”
@@
If the energy $E$ corresponds to an eigenstate that is localized, then its localized length $\Lambda=1/\gamma$ and the wave function of the eigenstate $\psi$ can be given by
$|\psi|\propto|\psi|_{\max}e^{-\gamma|j-j_{c}|},$ (S1)
where $|\psi|_{\max}$ is the maximum value of $|\psi|$ at the center site $j_{c}$ of localization. In Fig. S1(a) we show the fractal dimensions $\Gamma$ of all eigenstates in the complex plane for $\delta=1+i$, $h=1$, and plot the contours of the LE via $2\gamma(E)=\max\left\{\ln|\lambda a_{2}|+|h|,0\right\}$. The extension under non-Hermitian conditions originates from LE in the complex plane. The region with $\gamma=0$ forms one valley in the complex plane (inside the MR), and the eigenstate $\Gamma\rightarrow 1$ in this region is the extended state. The further away the eigenstate is from MR, the smaller $\Gamma$ is, and the more localized the corresponding eigenstate is. For eigenstates within MR, as shown in Fig. S1(b), the wavefunction can extend to all lattices. For localized states with $\gamma>0$, as shown in Fig. S1(c)(d), we present the distributions of the eigenstates for $\gamma=0.222$ and $\gamma=0.950$ over all lattices, respectively, and compare them with the fitted Eq. (S1). It is evident that Eq. (S1) effectively characterizes the localization features of the eigenstates, and the larger LE becomes, the more localized it is.
@@
$\eta\chi_{p}^{-1}(\omega_{0},0)e^{p\theta}u_{p}=G\sum_{r}e^{-\omega|p-r|}e^{r\theta}u_{r},$ (S8)
@@
$\eta\chi_{m}^{-1}(\omega,q\pi)v_{m}=G\sum_{r}e^{-\omega_{0}|m-r|}v_{r}.$ (S9)
In the last step we define $z_{k}=\sum_{m}e^{im(2\pi\alpha k+\theta)}v_{m}$. We multiply Eq. (S9) by $e^{im(2\pi\alpha j+\theta)}$ and sum over $m$ to obtain
@@
$H(\phi,K)=\begin{pmatrix}V(\phi/n_{2})&1&\dots&e^{iK}\\ 1&V(2\phi n_{1}/n_{2}+\phi/n_{2})&\dots&0\\ \vdots&&\ddots&\vdots\\ e^{-iK}&\dots&1&V(2\phi n_{1}(n_{2}-1)/n_{2})+\phi/n_{2})\\ \end{pmatrix}.$ (S12)
@@
It can be clearly seen that under the transformation of $\phi\leftrightarrow K$ ($\phi^{\prime}\leftrightarrow K$), the eigenvalue remains unchanged when $\lambda=1$ ($\lambda=-1$). Putting the two conditions together, we get the self-duality condition $|\lambda|=1$.
@@
By $\phi\leftrightarrow K$ and $\phi^{\prime}\leftrightarrow K$, we can obtain its self-duality point as $E=(\pm 2-\lambda)/b$ and this result is in agreement with Ref. SGaneshan2015 . In addition, this method works well in solving models that do not have self-duality.
@@
Further, we contrast this by Avila’s global theory. we can calculate LE from the transfer matrix
@@
$\gamma_{B}(E,\epsilon)=\epsilon+\left|\frac{Eb+i\lambda\pm\sqrt{(Eb+i\lambda)^{2}-4b^{2}}}{4}\right|$ (S21)
@@
$\gamma(E,0)=\mathrm{max}\left\{\ln\left|\frac{Eb+i\lambda\pm\sqrt{(Eb+i\lambda)^{2}-4b^{2}}}{2(1+\sqrt{1-b^{2}})}\right|,0\right\}.$ (S22)
@@
for $\gamma(E)=1$, where $C=[(1+\sqrt{1-b^{2}})^{2}-b^{2}]/(1+\sqrt{1-b^{2}})$. It is obviously an ellipse.